# Vizing’s Theorem in Near-Linear Time

### v1

@@ 2       Let $\alpha,\beta\in[\Delta+1]$ be the two least common colors in $\chi$
[…]
@@ 6.1 The Subroutine Prime-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 6.1). It then proceeds in iterations, where in each iteration it samples a u-fan $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$ uniformly at random and attempts to prime $\operatorname*{\boldsymbol{f}}$ with the colors $\alpha$ and $\beta$. The subroutine maintains a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\{\alpha,\beta\}$-primed u-fans. The subroutine performs iterations until $|\Phi|=\Omega(\lambda/\Delta)$, after which we proceed to call the subroutine Activate-U-Fans to extend the coloring $\chi$ to edges in the u-fans in $\Phi$. The pseudocode in Algorithm 6 gives a formal description of the subroutine Prime-U-Fans.
1 $\Phi\leftarrow\varnothing$ and $\lambda\leftarrow|\mathcal{U}|$
@@ 16                   Add the u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\operatorname*{\mathcal{U}}$ and $\Phi$
Analysis of Prime-U-Fans
We say that an iteration of (the while loop in) Prime-U-Fans is successful if it adds a u-fan to the set $\Phi$. Otherwise, we say that the iteration fails. Note that the algorithm repeatedly performs iterations until it has performed at least $\lambda/(48\Delta)$ successful iterations. The following lemmas summarize the main properties of the subroutine Prime-U-Fans.
@@
Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and $y$ [^10] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths. Since the collection of u-fans $\mathcal{U}$ is separable, we know that at most one u-fan in $\mathcal{U}$ containing $x$ (resp. $y$) has $c^{\prime}$ assigned as the color missing at $x$ (resp. $y$). Thus, we must sample one of these $2$ u-fans for the algorithm to flip the path $P$. ∎
@@
where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 6.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 6.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 6.4. Applying Markov’s inequality, it follows that
@@
Since each u-fan in $\Phi$ has 3 vertices and at most $2\Delta$$\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths end at each of these vertices, we know that there are at most $6\Delta|\Phi|$ alternating paths that could be considered during some iteration that end at a u-fan in $\Phi$. By 6.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths. Thus, at least $|\mathcal{U}|-12\Delta|\Phi|$ of the u-fans in $\mathcal{U}$ are contained in $\mathcal{U}^{\star}$. It follows that
@@
where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 6.2. Since $\operatorname*{\boldsymbol{f}}$ is sampled uniformly at random from $\operatorname*{\mathcal{U}}$, it follows that $\Pr[\operatorname*{\boldsymbol{f}}\in\mathcal{U}^{\star}]\geq 1/2$. The lemma follows by applying a union bound. ∎
@@ 4       Activate the u-fan $\operatorname*{\boldsymbol{f}}$ by flipping the path $P$ or $P^{\prime}$
[…]
@@
6.3 Analysis of Color-U-Fans (Proof of Lemma 4.2)
Given a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans repeatedly calls Prime-U-Fans and Activate-U-Fans with the set $\mathcal{U}$ as described in Algorithm 5. It repeats this process for $\Delta/2$ iterations. It follows from Claims 6.2 and 6.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration. Thus, by Bernoulli’s inequality, we get that
@@ 7 Implementation and Data Structures
Implementing an Edge Coloring: Let $G=(V,E)$ and a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$. We implement an edge coloring $\chi:E\longrightarrow\mathcal{C}$ of $G$ using the following:
@@
  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$ [^11] We take this intersection with $[\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
@@
For each $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{g}}$ is contained at most $3$ times in $\psi$. Thus, the total space required to store the hashmap that implements $\psi$ is $O(|\operatorname*{\mathcal{U}}|)$. Since $\operatorname*{\mathcal{U}}$ is separable, the u-components in $\operatorname*{\mathcal{U}}$ are edge-disjoint, and thus $|\mathcal{U}|\leq m$. It follows that map $\psi$ can be stored with $O(m)$ space. For each $u\in V$, we can observe that $|C_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)$ since at most $\deg_{G}(u)$ many u-components in $\operatorname*{\mathcal{U}}$ contain the vertex $u$, and $|\overline{C}_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)+1$ since $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\subseteq[\deg_{G}(u)+1]$. Thus, the total space required to store the sets $\{C_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ and $\{\overline{C}_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ is $O(m)$.
@@
[…]
Acknowledgements
Part of this work was conducted while the first and second named authors were visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
Appendix A Logarithmic Factors and Small Values of $\Delta$
The main contribution and the primary message of our work is presenting the first near-linear time algorithm for $(\Delta+1)$-edge coloring and polynomial time factor improvements over prior work. As such, in the presentation of our algorithm and its analysis, we made no attempt to optimize the logarithmic factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability.
Nevertheless, we can indeed optimize our algorithm to achieve a runtime of only $O(m\log{\Delta})$ time (with high probability), which matches the classical bound of [CH82, COS01] for $\Delta$-coloring bipartite graphs (although that algorithm is deterministic). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case; our result improves prior $O(m\Delta^{4}\log\Delta)$ bounds of [BD24] exponentially in terms of dependence on $\Delta$.
Theorem A.1.
There is a randomized algorithm that, given any simple undirected graph $G$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{\Delta})$ time with high probability.
@@ Appendix A Logarithmic Factors and Small Values of $\Delta$
    In Lemma 4.1, we are losing an $O(\log{n})$ factor in the subroutine Reduce-U-Edges while turning u-edges into u-fans or coloring them. Specifically, in the proof of Lemma 5.13, we lose an additional $O(\log{n})$ factor when handling all u-edges (see the Harmonic series in the proof).
@@
    Our application of Lemmas 4.1 and 4.2 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 4.3).
  4. 4.
@@
All in all, these factors lead to our $O(m\log^{3}{n})$ time algorithm (note that the losses in Steps 1 and 2 do not compound as we run these algorithms sequentially and thus pay an additive factor of $O(\log{n})$ between the two of them, not a multiplicative one).
We now show how these log factors can be removed or replaced by $O(\log{\Delta})$ in the following. Throughout, we continue our assumption in Section 3.2 that the high probability events of the hash maps used in the algorithm happen and we ignore their randomness (e.g., if an algorithm is deterministic modulo use of hash maps, we simply refer to it as deterministic here).
Step 1: The following lemma shows the $O(\log{n})$-term of Lemma 4.1 can be easily circumvented.
Lemma A.2 (A slight modification of Lemma 4.1).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, does one of the following in $O(m+\lambda\cdot\Delta)$ time deterministically:
  1. 1.
    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
  2. 2.
    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans.
Proof Sketch.
The $O(\log{n})$ term of the runtime in the proof of Lemma 4.1 comes from the Harmonic series calculation in the proof of Lemma 5.13. Specifically, for any color $\alpha$, assuming we start with $\lambda_{\alpha}$ $\alpha$-primed u-edges, by 5.12, the runtime needed for “handling” the $i^{th}$ such edge is $O((m_{\alpha}+\lambda_{\alpha})/i)$ time (by handling, we mean either coloring this edge or turning it into a u-fan). In the proof of Lemma 4.1, using 5.11, this results in a total runtime of
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log\lambda_{\alpha})=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log{n}).$
However, a simple observation is that we do not need to handle all the $\lambda_{\alpha}$ $\alpha$-primed u-edges given that our final goal is to extend the coloring or create u-fans out of only a constant fraction of the initial $\lambda$ uncolored edges. Specifically, suppose our goal is to reduce the number of $\alpha$-primed u-edges to some $\gamma\cdot\lambda_{\alpha}$ for any arbitrary constant $\gamma\in(0,1)$; then, the runtime will be
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=\gamma\cdot\lambda_{\alpha}}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O\left(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\cdot\log{\left(\frac{\lambda_{\alpha}}{\gamma\cdot\lambda_{\alpha}}\right)}\right)=O(\Delta\cdot\lambda_{\alpha}+m_{\alpha}),$
given that $\gamma$ is a constant. As such, the time it takes to handle any arbitrary large constant fraction of u-edges will be only $O(m)$ (by summing over choices of $\alpha\in[\Delta+1]$). The rest of the proof follows exactly as in Lemma 4.1 by simply removing the small constant fraction of u-edges not handled from the graph for the rest of the proof. ∎
Step 2: We remove the extra $O(\log{n})$-term of Lemma 4.2 in a similar way although by focusing on the expected runtime of the algorithm (plus a crucial tail inequality). In the last step, we show how to recover the final result even from this weaker guarantee.
Lemma A.3 (A slight modification of Lemma 4.2).
@@
Proof Sketch.
The runtime for coloring each single edge in Lemma 4.2 is $O(m/\lambda)$ by Lemma 6.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 6.5. We emphasize that this is in an average sense: the algorithm Color-U-Fans first uses Prime-U-Fans to prime $\Omega(\lambda/\Delta)$ u-fans in expected $O(m/\lambda)$ time per u-fan and thus $O((\lambda/\Delta)\cdot(m/\lambda))=O(m/\Delta)$ expected total time, and then deterministically colors them in $O(m/\Delta)$ time using Activate-U-Fans. It then repeats this process $\Delta/2$ times to color $\Omega(\lambda)$ edges, implying that $\mathbb{E}\left[T\right]=O(m)$ time.
@@
We note Lemma A.3 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability. For smaller values of $\lambda$ (which happens only as a corner case in the algorithm [^15] The only case in our algorithm where this tail bound cannot be replaced with a high probability bound is when $m/\Delta=o(\log{n})$ and at the same time $\Delta\log^{2}\Delta=\omega(m\log{\Delta})$, which means $m=o(n\log{n})$ and yet $\Delta=\omega(n/\log{n})$.), we need the specific tail inequality proven in the lemma instead.
Step 3: In the next step, we provide a more fine-grained version of Lemma 4.3 that will be used in the last step of the argument to (somewhat) bypass the $O(\log{n})$ factor loss of the original lemma.
Lemma A.4 (A slight modification of Lemma 4.3).
There exists an absolute constant $b>0$ such that the following holds. There is an algorithm given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
@@
The algorithm is verbatim as in Lemma 4.3 by replacing Lemma 4.1 with Lemma A.2 and Lemma 4.2 with Lemma A.3: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain. Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemmas A.2 and A.3 for a total of $O(\log{(\lambda/\lambda_{0})})$ times. Moreover, the $\Delta\lambda$ terms in the runtime of Lemma A.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total. This proves the expected bound on the runtime, i.e., $\mathbb{E}\left[T\right]\leq T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$.
As for the tail bound, for every application of Lemma A.3 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma A.3, we have,
@@
Pr[total $\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
@@
Step 4: Finally, we show how to put these lemmas together to $(\Delta+1)$ edge color the graph in $O(m\log{\Delta})$ randomized time, and conclude the proof of Theorem A.1.
Before getting to the main argument, we need the following specialized concentration inequality that we prove here (although, its formulation appears general enough to be of use elsewhere also). The proof follows a standard moment generating function argument.
Proposition A.5.

### v3

@@ 2   Let $\alpha,\beta\in[\Delta+1]$ be the two least common colors in $\chi$
3 $\textnormal{{Prime-U-Fans}}(\operatorname*{\mathcal{U}},\alpha,\beta)$
4 $\textnormal{{Activate-U-Fans}}(\operatorname*{\mathcal{U}},\alpha,\beta)$
@@ 8.1 The Subroutine Prime-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 8.1). It then proceeds in iterations, where in each iteration it samples a u-fan $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$ uniformly at random and attempts to prime $\operatorname*{\boldsymbol{f}}$ with the colors $\alpha$ and $\beta$. The subroutine maintains a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\{\alpha,\beta\}$-primed u-fans. The subroutine performs iterations until $|\Phi|=\Omega(\lambda/\Delta)$, after which we proceed to call the subroutine Activate-U-Fans to extend the coloring $\chi$ to edges in the u-fans in $\Phi$. The pseudocode in Algorithm 7 gives a formal description of the subroutine Prime-U-Fans.
[…]
@@ 16         Add the u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\operatorname*{\mathcal{U}}$ and $\Phi$
Algorithm 7 $\textnormal{{Prime-U-Fans}}(\mathcal{U})$
Analysis of Prime-U-Fans
We say that an iteration of (the while loop in) Prime-U-Fans is successful if it adds a u-fan to the set $\Phi$. Otherwise, we say that the iteration fails (see Figure 8 for an illustration). Note that the algorithm repeatedly performs iterations until it has performed at least $\lambda/(48\Delta)$ successful iterations. The following lemmas summarize the main properties of the subroutine Prime-U-Fans.
@@
Figure 8: In this picture, $\alpha$ is red, $\beta$ is orange, $\gamma$ is blue, and $\delta$ is green. $(u^{\prime},v^{\prime},w^{\prime})\in\Phi$ is an existing $\{\alpha,\beta\}$-primed u-fan in $\Phi$. If we flip the $\{\alpha,\gamma\}$-alternating path at $u$ and the $\{\beta,\delta\}$-alternating paths at $v$ and $w$, then we would destroy the property that $\beta\in\mathsf{miss}_{\chi}(v^{\prime})$.
@@
Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and yy. [^11] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths. Since the collection of u-fans $\mathcal{U}$ is separable, we know that at most one u-fan in $\mathcal{U}$ containing $x$ (resp. $y$) has $c^{\prime}$ assigned as the color missing at $x$ (resp. $y$). Thus, we must sample one of these $2$ u-fans for the algorithm to flip the path $P$. ∎
@@
where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 8.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 8.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 8.4. Applying Markov’s inequality, it follows that
@@
Since each u-fan in $\Phi$ has 3 vertices and at most $2\Delta$ $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths end at each of these vertices, we know that there are at most $6\Delta|\Phi|$ alternating paths that could be considered during some iteration that end at a u-fan in $\Phi$. By 8.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths (see Figure 9 for an illustration). Thus, at least $|\mathcal{U}|-12\Delta|\Phi|$ of the u-fans in $\mathcal{U}$ are contained in $\mathcal{U}^{\star}$. It follows that
@@
where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 8.2. Since $\operatorname*{\boldsymbol{f}}$ is sampled uniformly at random from $\operatorname*{\mathcal{U}}$, it follows that $\Pr[\operatorname*{\boldsymbol{f}}\in\mathcal{U}^{\star}]\geq 1/2$. The lemma follows by applying a union bound.
[…]
Figure 9: In this picture, $\alpha$ is red, $\beta$ is orange, and $\operatorname*{\boldsymbol{f}}^{\prime}=(u^{\prime},v^{\prime},w^{\prime},\alpha,\beta)\in\Phi$ is a u-fan, and we have drawn 99 different $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths starting from $\operatorname*{\boldsymbol{f}}^{\prime}$.
@@ 4   Activate the u-fan $\operatorname*{\boldsymbol{f}}$ by flipping the path $P$ or $P^{\prime}$
Algorithm 8 $\textnormal{{Activate-U-Fans}}(\mathcal{U},\alpha,\beta)$
@@
8.3 Analysis of Color-U-Fans: Proof of Lemma 6.3
Given a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans repeatedly calls Prime-U-Fans and Activate-U-Fans with the set $\mathcal{U}$ as described in Algorithm 6. It repeats this process for $\Delta/2$ iterations. It follows from Claims 8.2 and 8.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration. Thus, by Bernoulli’s inequality, we get that
@@ 9 Implementation and Data Structures
Implementing an Edge Coloring: Let $G=(V,E)$ be a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$. We implement an edge coloring $\chi:E\longrightarrow\mathcal{C}$ of $G$ using the following:
@@
  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$. [^12] We take this intersection with $\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
@@
For each $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{g}}$ is contained at most $3$ times in $\psi$. Thus, the total space required to store the hashmap that implements $\psi$ is $O(|\operatorname*{\mathcal{U}}|)$. Since $\operatorname*{\mathcal{U}}$ is separable, the u-components in $\operatorname*{\mathcal{U}}$ are edge-disjoint, and thus $|\mathcal{U}|\leq m$. It follows that the map $\psi$ can be stored with $O(m)$ space. For each $u\in V$, we can observe that $|C_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)$ since at most $\deg_{G}(u)$ many u-components in $\operatorname*{\mathcal{U}}$ contain the vertex $u$, and $|\overline{C}_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)+1$ since $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\subseteq[\deg_{G}(u)+1]$. Thus, the total space required to store the sets $\{C_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ and $\{\overline{C}_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ is $O(m)$.
@@
Implementing $\textsc{Find-Component}_{\operatorname*{\mathcal{U}}}(x,c)$: We make a query to $\psi$ by checking if there is an element $\psi_{x}(c)$. If no such element is contained in $\psi$, then return fail. Otherwise, return the unique u-component in the set $\psi_{x}(c)$.
Implementing $\textsc{Missing-Color}_{\operatorname*{\mathcal{U}}}(x)$: Return an arbitrary color from the set $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$.
10 The Final Algorithm: Proof of Theorem 1.1
Up until this point in the presentation of our algorithm and its analysis, we made no attempt in optimizing the logarithmic runtime factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability in Theorem 6.1. We now show that we can further optimize the algorithm and obtain an $O(m\log{n})$ time algorithm and conclude the proof of Theorem 1.1.
@@
    Our application of Lemmas 6.2 and 6.3 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 6.4).
@@
All in all, these factors led to the $O(m\log^{3}{n})$ bound of our algorithm in Theorem 6.1.
We now show how these log factors can be reduced to a single one. The two main ideas are: (1) relaxing the requirement of Lemma 6.3 so that its runtime holds in expectation, plus a suitable tail bound. Then, instead of maintaining high probability bound on each invocation of this lemma, we only bound the runtime of all invocations of this lemma together with high probability; (2) relaxing the Eulerian partition technique to color most of the graph recursively, instead of the entire graph, and then taking care of the remaining uncolored edges at the end.
We start by presenting a more fine-grained version of two of our main technical lemmas in the proof of Theorem 6.1 (Lemma 6.3 and Lemma 6.4; recall that Lemma 6.2 already runs in linear time deterministically), which correspond to part (1) above, and then use these to present part (2) of above and conclude the proof.
10.1 Fine-Grained Variants of  Lemma 6.3 and Lemma 6.4
We remove the $O(\log{n})$-term of Lemma 6.3 by focusing on the expected runtime of the algorithm (plus a crucial tail inequality). In the next part, we show how to recover the final result even from this weaker guarantee.
Lemma 10.1 (A slight modification of Lemma 6.3).
@@
Proof.
The amortized runtime for coloring each single edge in Lemma 6.3 is $O(m/\lambda)$ by Lemma 8.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 8.5. We emphasize that this is in an average sense: the algorithm Color-U-Fans first uses Prime-U-Fans to prime $\Omega(\lambda/\Delta)$ u-fans in expected $O(m/\lambda)$ time per u-fan and thus $O((\lambda/\Delta)\cdot(m/\lambda))=O(m/\Delta)$ expected total time, and then deterministically colors them in $O(m/\Delta)$ time using Activate-U-Fans. It then repeats this process $\Delta/2$ times to color $\Omega(\lambda)$ edges, implying that $\mathbb{E}\left[T\right]=O(m)$ time.
@@
We note Lemma 10.1 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability. For smaller values of $\lambda$ (which happens only as a corner case in the algorithm [^16] The only case in our algorithm where this tail bound cannot be replaced with a high probability bound is when $m/\Delta=o(\log{n})$ and at the same time $\Delta\log^{2}\Delta=\omega(m\log{\Delta})$, which means $m=o(n\log{n})$ and yet $\Delta=\omega(n/\log{n})$.), we need the specific tail inequality proven in the lemma instead.
Furthermore, we provide a similarly fine-grained version of Lemma 6.4 that will be used in the last step of the argument to bypass the $O(\log{n})$ factor loss of the original lemma.
Lemma 10.2 (A slight modification of Lemma 6.4).
There is an algorithm that given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
@@
The algorithm is verbatim as in Lemma 6.4 by replacing Lemma 6.3 with Lemma 10.1: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain. Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemma 6.2 and Lemma 10.1 for a total of $O(\log{(\lambda/\lambda_{0})})$ times. Moreover, the $\Delta\lambda$ terms in the runtime of Lemma 6.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total. This proves the expected bound on the runtime, i.e., $\mathbb{E}\left[T\right]\leq T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$.
As for the tail bound, in each application of Lemma 10.1 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma 10.1, we have,
@@
$\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
@@
Before moving on from this subsection, we mention the following specialized concentration inequality that we need in order to be able to exploit the tail bounds proven in Lemma 10.1 and Lemma 10.2. The proof follows a standard moment generating function argument and is postponed to Appendix D.
Proposition 10.3.
@@
10.2 Proof of Theorem 1.1
We are now ready to use the more fine-grained versions of our main technical lemmas to $(\Delta+1)$ edge color the graph in $O(m\log{n})$ randomized time, and conclude the proof of Theorem 1.1.
The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 6.1. Notice that the runtime of this algorithm is even $O(m\log{\Delta})$ and not $O(m\log{n})$ (the distinction at this point is irrelevant for us in proving Theorem 1.1, but we state the result this way so we can use it later in Appendix C as well).
Lemma 10.4.
There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but (exactly) $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.